### 14 Mar 2024 v1

# Parafermions with symmetry-protected non-Abelian statistics

Abstract
Non-Abelian anyons have garnered extensive attention for obeying exotic non-Abelian statistics and potential applications to fault-tolerant quantum computation. Although the prior research has predominantly focused on non-Abelian statistics without the necessity of symmetry protection, recent progresses have shown that symmetries can play essential roles and bring about a notion of the symmetry-protected non-Abelian (SPNA) statistics. In this work, we extend the concept of SPNA statistics to strongly-correlated systems which host parafermion zero modes (PZMs). This study involves a few fundamental results proved here. First, we unveil a generic unitary symmetry mechanism that protects PZMs from local couplings. Then, with this symmetry protection, the PZMs can be categorized into two nontrivial sectors, each maintaining its own parity conservation, even though the whole system cannot be dismantled into separate subsystems due to nonlinear interactions. Finally, by leveraging the parity conservation of each sector and the general properties of the effective braiding Hamiltonian, we prove rigorously that the PZMs intrinsically obey SPNA statistics. To further confirm the results, we derive the braiding matrix at a tri-junction. In addition, we propose a physical model that accommodates a pair of PZMs protected by mirror symmetry and satisfying the generic theory. This work shows a broad spectrum of strongly-correlated systems capable of hosting fractional SPNA quasiparticles and enriches our comprehension of fundamental quantum statistics linked to the symmetries that govern the exchange dynamics.
I Introduction
Quantum statistics of indistinguishable particles sets a fundamental principle in quantum mechanics Shankarbook , which classifies the elementary and quasi-particles into bosons, fermions, and anyons according to different basic rules upon their exchange Anyon1977 ; Wilczek1982 ; WilczekBook . The non-Abelian statistics adds to a captivating notion in the anyon statistics, which revolutionizes the traditional dichotomy of particles into bosons and fermions, triggering the paradigm shifts in both fundamental physics and potential applications Nayak1996 ; Ivanov2001 ; DasSarma2005 ; Nayak2008 ; Alicea2011 . Rather than a global statistical phase featuring the Abelian quantum statistics, the exchange of non-Abelian anyons is characterized by matrices which transform the quantum states of the many-body system. Such braiding statistics can be construed as a non-Abelian geometric phase defined in the degenerate ground states Berry1997 ; GeometricNRP2019 , and has driven the broad exploration of the exotic non-Abelian topological orders SternReview2010 . The quantum information can be encoded non-locally in non-Abelian anyons and manipulated by topological braiding with robustness to local noises, opening an elegant avenue for fault-tolerant quantum computation Kitaev2003 ; PachosBook ; Pachos2017 .
Among various non-Abelian anyons, the Majorana zero modes (MZMs) stand out due to the recent considerable experimental progress in the topological superconductors Mourik2012 ; MTDeng2012 ; Rokhinson2012 ; Shtrikman2012 ; JFJia2012 ; Marcus2013 ; JFJia2014 ; Yazdani2014 ; Marcus2015 ; Marcus2016 ; Molenkamp2016 ; Molenkamp2017 ; Qiao2017 ; HDing2018a ; HDing2018b ; Marcus2019 ; Molenkamp2019 ; Yazdani2019 ; Zhang2022 ; Microsoft2023 , where the Majorana modes emerge as quasiparticles identical to their antiparticles Kitaev2001 ; AliceaRPP ; Flensberg2012 ; Beenakker2013 ; Franz2015 ; Sato2017 . Two MZMs define a non-local complex fermion. The qubits encoded by these non-local fermions are robust against local perturbations Kitaev2001 . Braiding and measuring MZMs can implement Clifford gates, although these are not sufficient for universal quantum computation Gottesman1998 . From an algebraic standpoint, MZMs are considered as a special case within the broader category of $\mathbb{Z}_{N}$ parafermion zero modes (PZMs) with $N=2$ FradkinKadanoff ; Fendley2012 . Physically, the parafermions are fractionalized MZMs, emerging only in the strongly correlated systems SternPRX2012 ; MCheng2012 ; AliceaNC2013 ; Barkeshli2013 ; Barkeshli2014 ; Stern2014 ; Loss2014a ; Loss2014b ; Loss2014c ; Oreg2017general ; Oreg2017chiral ; Loss2019 ; Santos2020 ; Loss2020bilayer ; Oreg2020predicted ; Silva2022 . The braiding phases of PZMs also exhibit fractionalization compared to those of MZMs, manifesting a kind of non-Abelian fractional statistics. The parafermions can encode the nonlocal topological qudits FockPara , and braiding parafermion brings us closer to the computational universality than braiding MZMs AliceaNC2013 ; Loss2016 .
The non-Abelian statistics was initially introduced to singly existing MZMs which necessitate no symmetry protection Ivanov2001 ; Alicea2011 . The presence of symmetries can enrich the broad classes of topological superconductors SCZhang2009 ; TeoKane2010 ; Timm2010 ; Beenakker2011 ; KTLaw2012 ; Nagaosa2012 ; KaneMele2013 ; Berg2013 ; DLoss-a ; DLoss-b ; Nagaosa2014 ; magneticTSC ; Oreg2019 ; mirrorTSC ; Ueno2013 ; Sato2014crystalline ; C4TSC ; InversionTSC , in which a single topological defect may bind multiple zero modes and generate the ground state degeneracy. It was first shown in the time-reversal invariant topological superconductors that the Majorana Kramers pairs obey non-Abelian statistics due to the protection of time-reversal symmetry, rendering a new family of quantum statistics called the symmetry-protected non-Abelian (SPNA) statistics XJL2014PRX ; Gao2016 . This prediction shows an insight into the nontrivial roles of symmetries in the quantum statistics. The SPNA statistics has been further considered in various symmetry-protected topological superconductors Sato2014mirror ; JSH2022 ; YWuReview2023 ; NittaReview2024 , and is also the essential mechanism JSH2022 underlying the braiding statistics of Dirac zero modes Nitta2012dirac ; Loss2013fractional ; YWu2020a ; YWu2020b ; YWu2022 . The unitary and anti-unitary symmetries are fundamentally different in protecting the non-Abelian statistics. For the anti-unitary symmetric system, e.g. the time-reversal invariant topological superconductor, the non-Abelian statistics of Majorana Kramers pairs requires that not only the static topological superconductor satisfy the time-reversal symmetry, but also the braiding operation, which characterizes a dynamical evolution, satisfy time-reversal symmetry individually Gao2016 . This is because the time-reversal symmetry for the static superconductor cannot guarantee the absence of local coupling between two MZMs of each Kramers pair in the dynamical evolution localmixing1 ; localmixing2 ; Knapp2020 due to the profound dynamical symmetry-breaking of anti-unitary symmetries Gao2016 ; JasonAlicea2020. The dynamical symmetry breaking with coupling to environment opens up active discussions on the robustness of time-reversal invariant topological phases in the open systems Cooper2020 ; Zhai2021 ; Cai2021 . In contrast, the Majorana modes protected by unitary symmetries are intrinsic SPNA anyons without suffering the dynamical symmetry breaking JSH2022 . In particular, when the unitary symmetry is non-Abelian, the whole system cannot be block diagonalized into decoupled sectors corresponding to different symmetry eigenvalues, while the non-Abelian statistics is generically protected by the symmetry JSH2022 . The SPNA statistics elucidates that the quantum statistics of indistinguishable (quasi)-particles is essentially a dynamical rather than static fundamental feature. In additional to fundamental relevance, the symmetry-protected MZMs enrich new schemes for topological quantum computing JSH2022 ; OregReview2019 ; MCheng2021 ; TanakaMKP2022 ; LFu2022 .
Topological superconductors belong to the free-fermion topological phases XGWen2012 , for which the SPNA statistics has been so far proposed and thoroughly studied. Whether or not this concept can be generalized to strongly correlated systems, in particular those with PZMs, is an open question with fundamental importance, but confronts formidable challenges. First, for parafermions it was not even known how to generally introduce the symmetry protection. Further, due to strong interactions, systems accommodating PZMs cannot be dismantled into separate symmetry sectors, bringing the challenges for generic study. Finally, the braiding transformation of PZMs is nonlinear Loss2016 , and the characterization with symmetries can be even more complicated. Nevertheless, generalizing the SPNA statistics to strongly correlated systems could bring a new horizon in the fundamental physics of quantum statistics and promote intriguing applications to quantum computing.
In this work, we propose the theory of non-Abelian statistics for parafermions protected by unitary symmetry, and present a systematic study consisting of the profound results. We first unveil the generic symmetry mechanism for protecting the PZMs, as formulated with a unitary transformation on the anyon parities, with which the local couplings are prohibited. With this symmetry protection, we show that the PZMs can be categorized into two nontrivial symmetric sectors, each maintaining its own parity conservation, even though the whole system cannot be dismantled into separate subsystems due to nonlinear interactions. By harnessing the parity conservation of each sector and leveraging the fundamental properties of the effective braiding Hamiltonian, we prove rigorously that the PZMs intrinsically obey SPNA statistics, as captured by two independent braiding operations [see Fig. 1]. We further propose a physical model that accommodates a pair of PZMs protected by a mirror symmetry and satisfying the theory of SPNA statistics. The strong correlation effects are crucial for the model realization and affect the concrete form of the braiding matrix, but the SPNA statistics are generally obtained.
The rest of the article is organized as follows. In Sec. II, we provide a rigorous proof of the SPNA statistics of PZMs based on showing several theorems. Then in Sec. III, we derive the braiding matrix through a tri-junction, which further confirms the SPNA statistics of the parafermions. A concrete physical realization is proposed in Sec. IV through a strongly correlated nanowire model hosting PZMs with mirror symmetry. The conclusion and discussions are given in Sec. V, with the important future issues being commented.
Figure 1: Schematic diagrams for two pairs of PZMs residing at two sides of the system protected by a unitary symmetry. (a) The braiding of two pairs of PZMs reduces to (b) two individual sectors (the PZMs are redefined according to a sophisticated unitary transformation $\tilde{\alpha}_{i},\tilde{\beta}_{i}\rightarrow\alpha_{i},\beta_{i}$), each of which braids two PZMs independently. The wavy lines denote the interaction between two subsystems. These terms indicate that although the PZMs are categorized into two independent sectors, the subsystems are not independent ones.
II Non-Abelian braiding of PZMs with unitary symmetry protection
In this section, through a systematic study we prove rigorously that PZMs with unitary symmetry protection are intrinsic non-Abelian fractional quasiparticles. We first unveil the generic mechanism of unitary symmetry protection for the locally coexisting PZMs, for which all the local couplings are forbidden. This result leads to two key consequences. One is that the PZMs can be categorized into two separate sectors by redefining these modes from a nontrivial symmetric decomposition, yet the whole system cannot be decoupled into independent subsystems due to the strong interactions. The other is that in terms of the decomposed PZMs, the anyon parity of each symmetry sector is conserved. With the parity conservation and topological spin statistics theorem, we prove rigorously that the PZMs with unitary symmetry protection intrinsically obey the SPNA statistics.
II.1 Unitary symmetry protection
We first define the system of locally-coexisting PZMs and determine the generic form of the local coupling terms without symmetry protection. Consider two pairs of $\mathbb{Z}_{N}$ PZMs localized at two edges (domain walls) of the system with the commutation relations
$\tilde{\alpha}_{i}\tilde{\beta}_{j}$ $=$ $\omega\tilde{\beta}_{j}\tilde{\alpha}_{i},$ (1)
$\tilde{X}_{i}\tilde{X}_{j}$ $=$ $\omega^{\text{sgn}(j-i)}\tilde{X}_{j}\tilde{X}_{i},\ \tilde{X}=\tilde{\alpha},\tilde{\beta},$ (2)
where $\omega=\exp(2\pi i/N)$, $\tilde{\alpha}_{i=1,2}$, $\tilde{\beta}_{j=1,2}$ are unitary operators that commute with the system’s Hamiltonian and satisfy $\tilde{\alpha}_{i}^{N}=\tilde{\beta}_{i}^{N}=1$, and $\tilde{\alpha}_{1(2)}$, $\tilde{\beta}_{1(2)}$ localize at the left (right) edge of the system. Without symmetry protection, the PZMs $\tilde{\alpha}$ and $\tilde{\beta}$ of different flavours may be coupled together by local Hamiltonian $H_{\text{c}}=\sum_{k,l}\xi_{k,l}\tilde{\alpha}_{1}^{k}\tilde{\beta}_{1}^{l}+\sum_{m,n}\eta_{m,n}\tilde{\alpha}_{2}^{m}\tilde{\beta}_{2}^{n}$, where $\xi_{k,l}^{*}=\omega^{kl}\xi_{N-k,N-l}$ and $\eta_{m,n}^{*}=\omega^{mn}\eta_{N-m,N-n}$ for hermitianity. Since the coupling terms must conserve the total parity defined by $\tilde{Q}_{\text{tot}}=\omega\tilde{\alpha}_{1}\tilde{\alpha}_{2}^{\dagger}\tilde{\beta}_{1}\tilde{\beta}_{2}^{\dagger}$, we can determine generally the concrete form of the local couplings
$H_{\text{c}}=\sum_{m=1}^{N-1}\xi_{m}\tilde{Q}_{1}^{m}+\sum_{n=1}^{N-1}\eta_{n}\tilde{Q}_{2}^{n},$ (3)
where $\xi_{m}=\xi_{N-m}^{*}$ and $\eta_{n}=\eta_{N-n}^{*}$ for hermitianity and $\tilde{Q}_{i}=\omega^{(N+1)/2}\tilde{\alpha}_{i}\tilde{\beta}_{i}^{\dagger}$ denotes the local $\mathbb{Z}_{N}$ parity. The Hamiltonian Eq. (3) describes the generic local coupling terms in the absence of symmetry protection other than the intrinsic total parity conservation. The presence of the these terms pushes the parafermion modes out of zero energy, and the role of symmetry protection is then to forbid these terms as detailed later.
We show that the generic unitary symmetry protection condition for a pair of PZMs can be expressed by the action of a unitary symmetry at the local parity operator. By definition, the most general symmetry protection of these PZMs is $[H_{\text{c}},S]\neq 0$, for $H_{\text{c}}$ in any form, where $S$ is a symmetry of the system. When $N=2$, $\tilde{\alpha}_{i}$ and $\tilde{\beta}_{i}$ are reduced to MZMs, then only two terms $i\xi_{1}\tilde{\alpha}_{1}\tilde{\beta}_{1}$ and $i\eta_{1}\tilde{\alpha}_{2}\tilde{\beta}_{2}$ survive, and they can be eliminated by a unitary symmetry $Si\tilde{\alpha}_{i}\tilde{\beta}_{i}S^{-1}=-i\tilde{\alpha}_{i}\tilde{\beta}_{i}$ JSH2022 . When $N>2$, there is more than one coupling term at a single edge, thus the symmetry protection condition is more complicate. To facilitate the derivation, we work in the bases of eigenstates of local parity operator $\tilde{Q}_{i}$. The generic condition $[H_{\text{c}},S]\neq 0$ implies that the unitary symmetry $S$ must be a cyclic permutation of the eigenstates with $N$ elements provided that $S|q\rangle$ is still a local parity eigenstate. In this case, there exists a unitary matrix $G$ such that $GSG^{-1}=e^{i\vartheta/N}V^{p}$, where $p$ and $N$ are coprime, $\vartheta$ is a phase factor and $V$ is the shift matrix $V|q\rangle=|q-1\rangle$ (the details are provided in Appendix A). This observation can be rewritten in the operator form
$SQ_{i}S^{-1}=\omega^{p}Q_{i},$ (4)
where $Q_{i}=G^{-1}\tilde{Q}_{i}G$ is the local parity operator of PZMs $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$ and $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. This condition Eq. (4) suggests that the symmetry can advance the new local parity $Q_{i}$ by $p$-increments, where $p$ and $N$ are coprime. For $N=2$, this means that the symmetry operation switches odd/even fermion parity to even/odd one.
II.2 Parity conservation for each sector
In this subsection, we prove that the generic condition for unitary symmetry protection Eq. (4) implies that the PZMs can be classified into two sectors, with the parity of each sector being conserved. We can show generically that the unitary symmetry $S$ operates on the PZMs as (see Appendix B for more details)
$S\alpha^{\prime}_{i}S^{-1}$ $=$ $\omega^{p_{1}}\alpha^{\prime}_{i}e^{i\sum_{n=1}^{N-1}\mu_{n}Q_{i}^{n}},$ (5)
$S\beta^{\prime}_{i}S^{-1}$ $=$ $\omega^{p_{2}}\beta^{\prime}_{i}e^{i\sum_{n=1}^{N-1}\mu_{n}Q_{i}^{n}},$ (6)
where $\mu_{n}^{*}=\mu_{N-n}$, $p_{1}$, $p_{2}$ are integers that satisfy $p_{1}-p_{2}=p$, and $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$, $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. We see that the symmetry operator couples $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$. By defining new PZMs via unitary transformation $\alpha_{i}=W\alpha^{\prime}_{i}W^{-1}$, $\beta_{i}=W\beta^{\prime}_{i}W^{-1}$, where $W=\exp\left({i\sum_{n=1}^{N-1}\nu_{n}Q_{i}^{n}}\right)$, $\nu_{n}=-\mu_{n}/[(\omega^{n}-1)(\omega^{pn}-1)]$, we find the symmetry protection condition can be rewritten in a decoupled form
$S\alpha_{i}S^{-1}=\omega^{p_{1}}\alpha_{i},\quad S\beta_{i}S^{-1}=\omega^{p_{2}}\beta_{i},$ (7)
which indicates that $\alpha_{i}$ and $\beta_{i}$ are eigenmodes of the symmetry. When $N=2$, the above procedure can be reduced to the Majorana case. For a pair of MZMs $\alpha^{\prime}_{i}$ and $\beta^{\prime}_{i}$, the symmetry acts as $S\alpha^{\prime}_{i}S^{-1}=\pm(\cos\phi\alpha^{\prime}_{i}+\sin\phi\beta^{\prime}_{i})$, $S\beta^{\prime}_{i}S^{-1}=\mp(-\sin\phi\alpha^{\prime}_{i}+\cos\phi\beta^{\prime}_{i})$, which give $\mu_{1}=\phi$. Then by setting $\nu_{1}=-\phi/4$, we have $\alpha_{i}=\cos\frac{\phi}{2}\alpha^{\prime}_{i}+\sin\frac{\phi}{2}\beta^{\prime}_{i}$, $\beta_{i}=-\sin\frac{\phi}{2}\alpha^{\prime}_{i}+\cos\frac{\phi}{2}\beta^{\prime}_{i}$, and it’s easy to verify that $S\alpha_{i}S^{-1}=\pm\alpha_{i}$, $S\beta_{i}S^{-1}=\mp\beta_{i}$. In other words, the MZMs $\alpha_{i}$ and $\beta_{i}$ are eigenstates corresponding to different eigenvalues ±$\pm 1$ of the unitary symmetry.
According to the decoupled form of unitary symmetry protection Eq. (7), the parity of each sector defined by $Q_{\alpha}=\omega^{(N+1)/2}\alpha^{\prime}_{1}{\alpha^{\prime}}_{2}^{\dagger}=\omega^{(N+1)/2}\alpha_{1}\alpha_{2}^{\dagger}$, $Q_{\beta}=\omega^{(N+1)/2}\beta^{\prime}_{1}{\beta^{\prime}}_{2}^{\dagger}=\omega^{(N+1)/2}\beta_{1}\beta_{2}^{\dagger}$ should be conserved. To see this, consider the generic non-local four-body interaction term $H_{I}=\xi_{k,l,m,n}\alpha_{1}^{k}\alpha_{2}^{l}\beta_{1}^{m}\beta_{2}^{n}+h.c.$, where $k+l+m+n=0\text{ mod }N$ due to the total parity conservation. The symmetry protection condition Eq. (7) further requires $k+l=m+n=0\text{ mod }N$, then the interaction term reads $H_{I}=\xi_{k,l,m,n}Q_{\alpha}^{k}Q_{\beta}^{m}+h.c.$, which commutes with the parity operator of each sector. Note that the interaction term couples $\alpha_{i}$-sector and $\beta_{i}$-sector when $k\neq 0$, $m\neq 0$, we remark that the parity conservation of each sector does not imply the decomposition of the whole system into two distinct subsystems.
II.3 Effective braiding Hamiltonian
Figure 2: Topological spin of Abelian anyon $\psi_{2,1}$. (a) $\psi_{2,1}$ is a composite of two anyons of the type $\psi_{1,0}$ (denoted by the red solid lines) and one anyon of the type $\psi_{0,1}$ (denoted by the blue dashed lines). The topological spin is defined as the phase caused by $2\pi$ rotation of this composite object relative to the rest of the system. The rotation in (a) is topologically equivalent to the multiple crossings in (b). The red-red, blue-blue and red-blue crossings give phase factors of $\phi_{\alpha}$, $\phi_{\beta}$ and $\varphi/2$, thus we obtain the topological spin $\theta_{2,1}=\exp[i(4\phi_{\alpha}+\phi_{\beta}+2\varphi)]$. More generally, the topological spin of anyon $\psi_{q_{\alpha},q_{\beta}}$ is $\theta_{q_{\alpha},q_{\beta}}=\exp[i(q_{\alpha}^{2}\phi_{\alpha}+q_{\beta}^{2}\phi_{\beta}+q_{\alpha}q_{\beta}\varphi)]$.
[…]
$H_{E}=\sum_{m=1}^{N-1}a_{m}Q_{\alpha}^{m}+\sum_{n=1}^{N-1}b_{n}Q_{\beta}^{n}+\sum_{k,l=1}^{N-1}c_{k,l}Q_{\alpha}^{k}Q_{\beta}^{l},$ (8)
where $a_{m}^{*}=a_{N-m}$, $b_{n}^{*}=b_{N-n}$, $c_{k,l}^{*}=c_{N-k,N-l}$ for hermitianity. The first and second terms in Eq. (8) denote the exchange Hamiltonian of two edges. The third term in $H_{E}$ couples $\alpha_{i}$-sector and $\beta_{i}$-sector and indicates that, generally speaking, time-periodic dynamics involving two pairs of PZMs cannot be divided into two independent processes. Thus to prove the SPNA statistics, we then explore certain conditions that are specific to braiding process to eliminate the third term in Eq. (8).
The braiding of two pairs of PZMs should satisfy anyon theory and spin-statistics theorem. The parity conservation of each sector suggests the following fusion rules of unitary symmetry-protected PZMs
$F\times F$ $=$ $\sum_{q_{\alpha},q_{\beta}=0}^{N-1}\psi_{q_{\alpha},q_{\beta}},$ (9)
$\psi_{q_{\alpha},q_{\beta}}\times\psi_{q^{\prime}_{\alpha},q^{\prime}_{\beta}}$ $=$ $\psi_{q_{\alpha}\oplus q^{\prime}_{\alpha},q_{\beta}\oplus q^{\prime}_{\beta}},$ (10)
where $a\oplus b=(a+b)\text{ mod }N$, $F$ denotes a domain wall with a pair of PZMs, or equivalently, a single zero-energy Fock parafermion mode FockPara and $\psi_{q_{\alpha},q_{\beta}}$ denotes the Abelian anyon with $q_{\alpha(\beta)}$ quasiparticles of $\alpha_{i}(\beta_{i})$-sector. Since the Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$ is a composite object, its topological spin $\theta_{q_{\alpha},q_{\beta}}$ can be deduced from those of $\psi_{1,0}$, $\psi_{0,1}$ and $\psi_{1,1}$. Suppose that the topological spins of $\psi_{1,0}$, $\psi_{0,1}$, $\psi_{1,1}$ are $\theta_{1,0}=\exp(i\phi_{\alpha})$, $\theta_{0,1}=\exp(i\phi_{\beta})$, $\theta_{1,1}=\exp[i(\phi_{\alpha}+\phi_{\beta}+\varphi)]$ separatedly, then the general expression for topological spin is $\theta_{q_{\alpha},q_{\beta}}=\exp[i(q_{\alpha}^{2}\phi_{\alpha}+q_{\beta}^{2}\phi_{\beta}+q_{\alpha}q_{\beta}\varphi)]$ as shown in Fig. 2 where $\theta_{2,1}$ is depicted as an example. The $\mathbb{Z}_{N}\times\mathbb{Z}_{N}$ structure of $\theta_{q_{\alpha},q_{\beta}}$ enforced by the fusion rules further requires $\theta_{q_{\alpha}+N,q_{\beta}}=\theta_{q_{\alpha},q_{\beta}}$ and $\theta_{q_{\alpha},q_{\beta}+N}=\theta_{q_{\alpha},q_{\beta}}$, which give
$\theta_{q_{\alpha},q_{\beta}}=\exp\left[i\frac{\pi}{N}\left(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w\right)\right],$ (11)
where $u$, $v$, $w$ are integer numbers. Label the eigenstate of the system $|q_{\alpha},q_{\beta}\rangle$ by the presence of Abelian anyon $\psi_{q_{\alpha},q_{\beta}}$, then the braiding operator $U(T)$ takes a matrix form with only diagonal matrix element $\left[U(T)\right]_{q_{\alpha},q_{\beta}}=\langle q_{\alpha},q_{\beta}|U(T)|q_{\alpha},q_{\beta}\rangle$. From the spin-statistics theorem $\theta_{q_{\alpha},q_{\beta}}=\theta_{F}^{2}\left[U(T)\right]_{q_{\alpha},q_{\beta}}^{2}$ Kitaev2006 , where $\theta_{F}$ denotes the topological spin of a Fock parafermion mode, we find
$\left[U(T)\right]_{q_{\alpha},q_{\beta}}=(-1)^{f(q_{\alpha},q_{\beta})}e^{i\pi(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)/2N},$ (12)
where ${f(q_{\alpha},q_{\beta})}$ is an integer-valued function. This formula indicates that the matrix elements of $U(T)$ can only take discrete values and $\left[U(T)\right]^{4N}=1$. From Eq (12), we can also read the diagonal matrix element of the effective braiding Hamiltonian $\langle q_{\alpha},q_{\beta}|H_{E}|q_{\alpha},q_{\beta}\rangle=(q_{\alpha}^{2}u+q_{\beta}^{2}v+2q_{\alpha}q_{\beta}w)[\pi/(2NT)]+(\pi/T)f(q_{\alpha},q_{\beta})$, thus the originally continuous coefficients $a_{m}$, $b_{n}$, $c_{k,l}$ in Eq. (8) should be similarly restricted to discrete values.
The symmetric effective braiding Hamiltonian Eq. (8) and the property $\left[U(T)\right]^{4N}=1$ are sufficient to the proof to unitary SPNA statistics of PZMs. The property $\left[U(T)\right]^{4N}=1$ or equivalently the discrete values of the matrix elements of $U(T)$ imply that the coefficients $(a_{m},b_{n},c_{k,l})$ in the effective braiding Hamiltonian must be discrete values which cannot change continuously. Thus the specific values of the coefficients can be determined by a special case as long as the Hamiltonian in the special case is topologically equivalent to that in the generic case. Futhermore, in a special case where the two sectors of PZMs $\alpha_{1,2}$ and $\beta_{1,2}$ belong to two fully decoupled subsystem copies, we must have $c_{k,l}=0$ for all the $k$, $l$. Thus the braiding operator should also be decoupled for the generic case where the whole system cannot be decoupled into two independent subsystems. In this case, the PZMs are braided independently in each sector. Then the braiding operator reads
$U(T)=\frac{1}{N}\left(\sum_{m=0}^{N-1}\lambda_{\alpha,m}Q_{\alpha}^{m}\right)\otimes\left(\sum_{n=0}^{N-1}\lambda_{\beta,n}Q_{\beta}^{n}\right),$ (13)
where $\lambda_{\alpha(\beta),m}=\omega^{m(m+N+2r_{\alpha(\beta)})/2}$ with $r_{\alpha(\beta)}=0,\cdots,N-1$. This formula is the main result of this work and gives the first example of SPNA statistics in correlated systems. Our proof only relys on the unitary symmetry protection and the general property of braiding process. When $N=2$, the Eq. (13) is reduced to the braiding operator of two symmetry-protected Majorana doublets XJL2014PRX ; Gao2016 ; JSH2022 or two Dirac zero modes Loss2013fractional ; YWu2020a ; YWu2020b ; YWu2022 .
III Braiding matrix in a tri-junction
Figure 3: Braiding process in a tri-junction. (a) Four pairs of PZMs $\alpha_{1,\cdots,4}$ and $\beta_{1,\cdots,4}$ constitude a two-$N$-qudit system (with definite total parity of each sector) and $\alpha_{5,6}$, $\beta_{5,6}$ serve as ancilla modes to form a tri-junction and braid PZMs in the qudits. The system is initialized by coupling $\alpha_{5,6}$, $\beta_{5,6}$ together. (b1-b3) During the braiding of $\alpha_{1}(\beta_{1})$ and $\alpha_{2}(\beta_{2})$, the PZMs $\alpha_{3,4}$, $\beta_{3,4}$ are decoupled from the rest PZMs. The ancilla modes $\alpha_{6}(\beta_{6})$ are coupled with $\alpha_{5}(\beta_{5})$, $\alpha_{1}(\beta_{1})$, $\alpha_{2}(\beta_{2})$, $\alpha_{5}(\beta_{5})$ in succession. The PZMs are effectively transported according to the arrows. The groundstate degeneracy is kept as $N^{2}$-fold in the process. (c1-c3) Time dependence of the coupling strength. $t_{i}^{\text{II}}=t_{f}^{\text{I}}$ and $t_{i}^{\text{III}}=t_{f}^{\text{II}}$. From $t_{i}^{\text{I}}$ to $t_{f}^{\text{III}}$, the Hamiltonian is changed back to its original form while the PZMs $\alpha_{1}$, $\beta_{1}$ and $\alpha_{2}$, $\beta_{2}$ are exchanged adiabatically.
We can also establish the SPNA statistics by explicitly deriving the non-Abelian geometric phases resulting from the manipulation of PZMs within a tri-junction configuration. The exchange of PZMs is accomplished through a three-point turn in the tri-junction, with the groundstate degeneracy remaining invariant throughout the adiabatic exchange process. In the following subsections, we expound upon the braiding dynamics within the tri-junction and delve into the characteristics of degenerate groundstates of the system. Subsequently, we derive the recursion relation for non-Abelian Berry phases accumulated during the braiding process. Finally, we demonstrate that the Berry phases necessarily adopt a decoupled form, as dictated by the adiabatic condition.
III.1 Tri-junction and groundstate subspace
The braiding process consists of three subsequent stages of adiabatic time evolution. Consider six pairs of symmetry-protected PZMs $\alpha_{1,\cdots,6}$ and $\beta_{1,\cdots,6}$ as shown in Fig. 3(a) and suppose that we are going to braid $\alpha_{1}$, $\beta_{1}$ and $\alpha_{2}$, $\beta_{2}$. Before the braiding, we initialize the system by coupling $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ together, then only $\alpha_{1,\cdots,4}$ and $\beta_{1,\cdots,4}$ remain zero-energy. Since the total parity for each sector is conserved for symmetry-protected parafermionic systems during the braiding, if the parity for $\alpha(\beta)$-sector is fixed, then the groundstates after initialization of this system is $N\times N$-fold degenerate. At the first stage of braiding process, the coupling between $\alpha_{1}$, $\beta_{1}$ and $\alpha_{6}$, $\beta_{6}$ is turned on adiabatically, and the coupling between $\alpha_{5}$, $\beta_{5}$ and $\alpha_{6}$, $\beta_{6}$ is turned off adiabatically at the same time. The coupling Hamiltonian for this stage is written as $H_{\text{I}}(t)=\eta_{16}(t)H_{16}+\eta_{56}(t)H_{56}$, where $\eta_{16}(t)$, $\eta_{56}(t)$ are depicted in Fig. 3(c1) and $H_{ij}$ denotes the coupling term between $\alpha_{i}$, $\beta_{i}$ and $\alpha_{j}$, $\beta_{j}$, whose concrete form will be specified later. The second and third stages of the braiding process are governing by similar Hamiltonians $H_{\text{II}}(t)=\eta_{26}(t)H_{26}+\eta_{16}(t)H_{16}$ and $H_{\text{III}}(t)=\eta_{56}(t)H_{56}+\eta_{26}(t)H_{26}$ as depicted in Fig 3. After all the three stages, the Hamiltonian of the whole system is returned to the original form after initialization and this braiding process is finished. During the braiding, there are $N\times N$ degenerate groundstates, and since $\alpha_{3}$, $\beta_{3}$ and $\alpha_{4}$, $\beta_{4}$ are zero-modes at each stage, the instantaneous goundstates can always be labelled as the parity eigenstates of $\alpha_{3}$, $\alpha_{4}$ and $\beta_{3}$, $\beta_{4}$
$Q_{X,34}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t)\rangle=\omega^{q^{\prime}_{X}}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t)\rangle,$ (14)
where $Q_{X,34}=\omega^{(N+1)/2}X_{3}X_{4}^{\dagger}$ with $X=\alpha,\beta$, $q^{\prime}_{X}$ is $q_{X,34}$ for short. In these bases, the effect of braiding is manifested in the Berry phase of the groundstate $|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{f})\rangle=e^{i\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{i})\rangle$, where $t_{i(f)}$ denotes the initial (final) time of the braiding dynamics.
III.2 Recursion relation for Berry phases
[…]
$Q_{\alpha,23}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t)\rangle$ $=$ $e^{i\delta_{\alpha}^{\text{I}}}|\psi(q^{\prime}_{\alpha}+1,q^{\prime}_{\beta},t)\rangle,$ (15)
$Q_{\beta,23}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t)\rangle$ $=$ $e^{i\delta_{\beta}^{\text{I}}}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta}+1,t)\rangle,$ (16)
where $\delta_{\alpha(\beta)}^{\text{I}}$ is short for $\delta_{\alpha(\beta)}^{\text{I}}(q^{\prime}_{\alpha},q^{\prime}_{\beta},t)$. And we have $U_{\text{I}}Q_{X,23}|\psi(t^{\text{I}}_{i})\rangle=e^{i\chi_{\text{I}}}Q_{X,23}|\psi(t^{\text{I}}_{f})\rangle$ with $X=\alpha$, $\beta$ since $U_{\text{I}}$ and $Q_{X,23}$ commute. It follows that $\chi_{\text{I}}(q^{\prime}_{X}+1)=\chi_{\text{I}}(q^{\prime}_{X})+\delta_{X}^{\text{I}}(t_{f}^{\text{I}})-\delta_{X}^{\text{I}}(t_{i}^{\text{I}})$. Similar identities can be obtained for the second (third) stage by replacing the symmetry operators $Q_{X,23}$ by $Q_{X,35(13)}$. By defining $\delta_{X,i(f)}=\delta_{X}^{\text{I}}(t_{i(f)}^{\text{I}})+\delta_{X}^{\text{II}}(t_{i(f)}^{\text{II}})+\delta_{X}^{\text{III}}(t_{i(f)}^{\text{III}})$, we obtain
$\chi(q^{\prime}_{\alpha}+1,q^{\prime}_{\beta})$ $=$ $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})+\delta_{\alpha,f}-\delta_{\alpha,i},$ (17)
$\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta}+1)$ $=$ $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})+\delta_{\beta,f}-\delta_{\beta,i}.$ (18)
From the above recursion relations, we see that if the phase factor $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ can be divided into two independent parts $\delta_{X,i(f)}(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\delta^{\alpha}_{X,i(f)}(q^{\prime}_{\alpha})+\delta^{\beta}_{X,i(f)}(q^{\prime}_{\beta})$, where $\delta^{\alpha}_{X,i(f)}(\delta^{\beta}_{X,i(f)})$ only relies on a single parity $q^{\prime}_{\alpha}(q^{\prime}_{\beta})$, then the total Berry phase $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ is also decoupled.
III.3 Decoupled braiding matrix
The adiabatic condition implies the decoupled forms of the phase factor $\delta_{X,i(f)}$ and the non-Abelian Berry phases. To see this, we explore the concrete form of the coupling Hamiltonian $H_{ij}$. Most generically,
$H_{ij}=H_{ij}^{\alpha}+H_{ij}^{\beta}+H_{ij}^{\text{int}},$ (19)
[…]
$Q_{X,56}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{i}^{\text{I}})\rangle$ $=$ $\omega^{q_{X,56}}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{i}^{\text{I}})\rangle,$ (20)
$Q_{X,16}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{f}^{\text{I}})\rangle$ $=$ $\omega^{q_{X,16}}|\psi(q^{\prime}_{\alpha},q^{\prime}_{\beta},t_{f}^{\text{I}})\rangle.$ (21)
Certain Hamiltonians like $H_{ij}^{\alpha}=H_{ij}^{\beta}=0$, $H_{ij}^{\text{int}}=c^{\prime}_{1,1}Q_{\alpha,ij}Q_{\beta,ij}+h.c.$ are forbidden since their groundstates are degenerate. Then the physical coupling Hamiltonian must be continuously connected to the simple one $H^{\prime}_{ij}=a^{\prime}_{1}Q_{\alpha,ij}+b^{\prime}_{1}Q_{\beta,ij}+h.c.$, whose groundstate is non-degenerate. The Hamiltonian $H^{\prime}_{ij}$ takes a decoupled form, as a result, the braiding matrix can be divided into two independent matrices for the two sectors of PZMs
$\gamma(q^{\prime}_{\alpha},q^{\prime}_{\beta})=\frac{\pi}{N}\left[(q^{\prime}_{\alpha}-k_{\alpha})^{2}+(q^{\prime}_{\beta}-k_{\beta})^{2}\right],$ (22)
where $k_{\alpha(\beta)}$ is integer. This result can also be obtained by effectively regarding the braiding process as successive forced measurements Bonderson2008 . The Eqs. (20) and (21) indicate that during each stage, the parity of each sector can be measured independently, and thus the braiding phases can be divided into two independent parts. Switch the focus from parity eigenstates to zero-mode operators, we find the non-Abelian Berry phases $\chi(q^{\prime}_{\alpha},q^{\prime}_{\beta})$ correspond to the following transformation of PZM operators
$\begin{split}\alpha_{1}\to s_{\alpha}\alpha_{2},&&\alpha_{2}\to\omega s_{\alpha}\alpha_{1}^{\dagger}\alpha_{2}^{2},\\ \beta_{1}\to s_{\beta}\beta_{2},&&\beta_{2}\to\omega s_{\beta}\beta_{1}^{\dagger}\beta_{2}^{2},\end{split}$ (23)
where $s_{\alpha(\beta)}=\omega^{k_{\alpha(\beta)}+N/2}$. This transformation is also consistent with Eq. (13) with $k_{\alpha(\beta)}=-r_{\alpha(\beta)}-N/2$. The above result shows that each PZM in one pair sees only one PZM rather than both parafermion modes in another pair during the braiding process, manifesting a central feature of the SPNA statistics for parafermions.
IV Symmetry-protected PZMs in interacting quantum wires
We provide a nanowire model hosting a pair of PZMs at each edge protected by mirror symmetry. In this model, the PZMs are naturally classified into two sectors with conserved parity. However, symmetry-preserving interactions can mix the PZMs of the two sectors without ruining non-Abelian statistics, which serves as a non-trivial demonstration of our generic theory.
Figure 4: Mechanism for the mirror symmetric model and local mixing under perturbation. (a) Two spin-orbit coupled nanowires with opposite spin-orbit couplings are placed in between two conventional superconductors with $\pi$-phase difference. The nanowires are coupled by tunneling $M$. Pairs of PZMs are localized at the ends of the system. (b) The band structure of the nanowires when $M=\Delta=0$. (c) At fractional filling $\nu=1/3$, pairing terms (blue solid lines) and backscattering terms (red dashed lines) assisted by interactions are dominant, giving rise to $\mathbb{Z}_{6}$ PZMs. (d) Dimerized lattice model hosting symmetry-protected PZMs. The dashed line and circle denote the symmetric perturbation.
IV.1 Model with mirror symmetry
We start with a non-interacting model hosting pairs of symmetry-protected MZMs and then switch to the strongly-correlated cases. As depicted in Fig. 4(a), the model consists of two spin-orbit coupled nanowires proximited to two conventional superconductors with $\pi$-phase difference and the Bloch Hamiltonian reads
$\mathcal{H}(k)=(\varepsilon_{k}-\mu)\tau_{z}-\alpha k\eta_{z}\sigma_{z}+M\eta_{x}\tau_{z}+\Delta\eta_{z}\tau_{y}\sigma_{y},$ (24)
where $\eta_{i}$, $\tau_{i}$, $\sigma_{i}$ are Pauli matrices acting in left/right-side, particle-hole and spin subspaces. The spin-orbit couplings of the two nanowires are opposite and in the absence of superconductivity $\Delta$ and inter-wire tunneling $M$, the band structure is as shown in Fig. 4(b). This system respects time-reversal symmetry $\mathcal{T}\mathcal{H}(k)\mathcal{T}^{-1}=\mathcal{H}(-k)$ and particle-hole symmetry $\mathcal{P}\mathcal{H}(k)\mathcal{P}^{-1}=-\mathcal{H}(-k)$ with $\mathcal{T}=i\sigma_{y}\mathcal{K}$ and $\mathcal{P}=\tau_{x}\mathcal{K}$, and thus belongs to the symmetry class DIII. The system is also invariant under a unitary mirror symmetry $\mathcal{M}\mathcal{H}(k)\mathcal{M}^{-1}=\mathcal{H}(k)$ with $\mathcal{M}=\eta_{x}\sigma_{x}$, and we can block-diagonalize the Hamiltonian $\mathcal{H}(k)$ with respect to the two eigenvalues of $\mathcal{M}$. In the topological regime, both blocks host MZMs as edge states. The two MZMs are eigenstates of different eigenvalues of the symmetry operator $\mathcal{M}\gamma\mathcal{M}^{-1}=\gamma$, $\mathcal{M}\bar{\gamma}\mathcal{M}^{-1}=-\bar{\gamma}$, and thus are protected by the mirror symmetry $\mathcal{M}i\gamma\bar{\gamma}\mathcal{M}^{-1}=-i\gamma\bar{\gamma}$. To realize symmetry-protected PZMs beyond MZMs, we tune the chemical potential to fractional filling $\nu=1/(2n+1)$, and introduce non-commutative interaction terms to open different gaps. These gaps effectively form domain walls at the edges of the system, leading to the appearance of PZMs. This physical picture is made clear by adopting the following bosonization formalism. The spectra around Fermi points are linearized as $\psi_{\eta^{\prime}\sigma^{\prime}}=R_{\eta^{\prime}\sigma^{\prime}}e^{ik_{F}^{1\eta^{\prime}\sigma^{\prime}}x}+L_{\eta^{\prime}\sigma^{\prime}}e^{ik_{F}^{\bar{1}\eta^{\prime}\sigma^{\prime}}x}$, where $\psi_{\eta^{\prime}\sigma^{\prime}}$ denote the bases after block-diagonalization $\mathcal{M}\psi_{\eta^{\prime}\sigma^{\prime}}\mathcal{M}^{-1}=\eta^{\prime}\psi_{\eta^{\prime}\sigma^{\prime}}$ and $\eta^{\prime}=\pm 1$, $\sigma^{\prime}=\pm 1$. The non-commutative backscattering and pairing terms that open different gaps are written as
$H_{M}$ $=\tilde{M}\sum_{\eta^{\prime}}[R^{\dagger}_{\eta^{\prime}\bar{\eta}^{\prime}}L_{\eta^{\prime}\bar{\eta}^{\prime}}]^{n}R^{\dagger}_{\eta^{\prime}\bar{\eta}^{\prime}}L_{\eta^{\prime}\eta^{\prime}}[R^{\dagger}_{\eta^{\prime}\eta^{\prime}}L_{\eta^{\prime}\eta^{\prime}}]^{n},$ (25)
$H_{\Delta}$ $=\tilde{\Delta}\sum_{\eta^{\prime}}[R^{\dagger}_{\eta^{\prime}\eta^{\prime}}L_{\eta^{\prime}\eta^{\prime}}]^{n}R^{\dagger}_{\eta^{\prime}\eta^{\prime}}L^{\dagger}_{\eta^{\prime}\bar{\eta}^{\prime}}[R_{\eta^{\prime}\bar{\eta}^{\prime}}L^{\dagger}_{\eta^{\prime}\bar{\eta}^{\prime}}]^{n},$ (26)
where $\bar{\eta}^{\prime}$ is short for $-\eta^{\prime}$. In the spetical case $n=1$, the filling is $\nu=1/3$, and the processes that constitude $H_{M}$ and $H_{\Delta}$ are depicted in Fig. 4(c). These terms can be rewritten as $H_{M}=2\tilde{M}\sum_{\eta^{\prime}}\cos[\varphi_{1\eta^{\prime}\bar{\eta}^{\prime}}-\varphi_{\bar{1}\eta^{\prime}\eta^{\prime}}]$ and $H_{\Delta}=2\tilde{\Delta}\sum_{\eta^{\prime}}\cos[\varphi_{1\eta^{\prime}\eta^{\prime}}+\varphi_{\bar{1}\eta^{\prime}\bar{\eta}^{\prime}}]$ after bosonization procedure with $\varphi_{r\eta^{\prime}\sigma^{\prime}}=(n+1)\phi_{r\eta^{\prime}\sigma^{\prime}}-n\phi_{\bar{r}\eta^{\prime}\sigma^{\prime}}$ and $R_{\eta^{\prime}\sigma^{\prime}}=e^{i\phi_{1\eta^{\prime}\sigma^{\prime}}}$, $L_{\eta^{\prime}\sigma^{\prime}}=e^{i\phi_{\bar{1}\eta^{\prime}\sigma^{\prime}}}$. By applying the unfolding method Stern2014 ; Giamarchi , the gaps opened by the two terms form a domain wall at edge, where we identify a pair of $\mathbb{Z}_{2(2n+1)}$ PZMs
$\alpha$ $=$ $e^{i[(\varphi_{11\bar{1}}-\varphi_{\bar{1}11})+(\varphi_{111}+\varphi_{\bar{1}1\bar{1}})]/[2(2n+1)]},$ (27)
$\beta$ $=$ $e^{i[(\varphi_{1\bar{1}1}-\varphi_{\bar{1}\bar{1}\bar{1}})+(\varphi_{1\bar{1}\bar{1}}+\varphi_{\bar{1}\bar{1}1})]/[2(2n+1)]}.$ (28)
These PZMs are protected by the mirror symmetry since $\mathcal{M}\alpha\mathcal{M}^{-1}=\alpha$ and $\mathcal{M}\beta\mathcal{M}^{-1}=\omega\beta$, satisfying the generic theory. Since we have defined the PZMs in the eigenbasis of the mirror symmetry, they are naturally classified into two symmetry sectors with conserved parity.
IV.2 Local mixing under symmetric perturbation
As mentioned in the generic theory, the parity conservation of each sector does not imply the decomposition of the whole system into two distinct subsystems. To show that certain interactions can mix modes in two sectors without ruining non-Abelian statistics, we model the above system with a lattice Hamiltonian (see Fig. 4(d))
$H_{\text{latt}}=H_{\alpha}+H_{\beta}+H_{p},$ (29)
where $H_{X}=\sum_{i=1}^{N-1}JX_{2i}^{\dagger}X_{2i+1}+J^{*}X_{2i+1}^{\dagger}X_{2i}$ and $\alpha_{i}$, $\beta_{i}$ are parafermionic operators that satisfy $\mathcal{M}\alpha_{i}\mathcal{M}^{-1}=\alpha_{i}$, $\mathcal{M}\beta_{i}\mathcal{M}^{-1}=\omega\beta_{i}$. The third part $H_{p}$ includes certain symmetric perturbation terms $H_{p}=f\alpha_{1}^{\dagger}\alpha_{2}+f^{*}\alpha_{2}^{\dagger}\alpha_{1}+V\alpha_{2}^{\dagger}\alpha_{3}\beta_{2}^{\dagger}\beta_{3}+V^{*}\beta_{3}^{\dagger}\beta_{2}\alpha_{3}^{\dagger}\alpha_{2}$, where $f/J\ll 1$, $V/J\ll 1$, $f/V\sim\mathcal{O}(1)$ and $[H_{p},\mathcal{M}]=0$. In the absence of $H_{p}$, the lattice system is in a dimerized phase and $\alpha_{1,N}$, $\beta_{1,N}$ are exact PZMs. The terms in $H_{p}$ only perturb a single zero mode $\alpha_{1}$ and the new PZM in the presence of $H_{p}$ can be constructed as $\bar{\alpha}_{1}=\alpha_{1}+\sum_{i}m_{i}\hat{O}_{i}^{(1)}+\sum_{j}n_{j}\hat{O}_{j}^{(2)}+\cdots$, where $m_{i}\sim\mathcal{O}(f/J)$ and $n_{j}\sim\mathcal{O}(f^{2}/J^{2})$. By calculating the perturbation series order by order, we find $\tilde{\alpha}_{1}$ cannot be constructed solely by terms with $\alpha_{i}$ , and its dynamics is governed by the full Hamiltonian rather than just $H_{\alpha}$ (see Appendix C). If $\alpha_{i}$ and $\beta_{i}$ are reduced to Majorana operators $\gamma_{i}^{\alpha}$ and $\gamma_{i}^{\beta}$, the exact zero mode can be constructed explicitly
$\bar{\gamma}_{1}^{\alpha}=\gamma_{1}^{\alpha}+a\gamma_{3}^{\alpha}+ib\gamma_{2}^{\beta}\gamma_{3}^{\alpha}\gamma_{3}^{\beta}+ic\gamma_{1}^{\alpha}\gamma_{2}^{\beta}\gamma_{3}^{\beta},$ (30)
where $a\sim\mathcal{O}(f/J)$, $b\sim\mathcal{O}(f^{2}/J^{2})$ and $c\sim\mathcal{O}(f^{3}/J^{3})$. From this expression, we find the Majorana operators $\beta_{2,3}$ in $H_{\beta}$ are involved in the wavefunction of $\bar{\gamma}_{1}^{\alpha}$, which is an MZM in $\alpha$-sector since $\mathcal{M}\bar{\gamma}_{1}^{\alpha}\mathcal{M}^{-1}=\bar{\gamma}_{1}^{\alpha}$. This confirms that interactions modify the concrete form of the braiding but not affecting the SPNA statistics in general.
V Conclusions and Discussions
In summary, we have shown that the PZMs with unitary symmetry obey the symmetry-protected non-Abelian (SPNA) statistics, manifesting a new paradigm of quantum statistics in strongly correlated systems, and established a systematic theory. The fractional topological phases hosting unitary symmetry-protected PZMs are strongly correlated and generically cannot be characterized through decoupled symmetry sectors, in sharp contrast to the free-fermion topological superconductors with symmetry-protected MZMs. Nevertheless, we have shown that the PZMs protected by unitary symmetry can always be classified into nontrivial symmetry sectors, with each sector preserving anyon parity individually in the dynamical evolutions. This profound symmetry protection mechanism determines the generic properties of the effective braiding Hamiltonian and, together with the anyon spin-statistics theorem, further leads to the SPNA statistics of PZMs. Finally, we have proposed a concrete physical model based on interacting nanowires to realized the PZMs with mirror symmetry, which satisfies the generic theory of the SPNA statistics.
This study advances the SPNA statistics from free-fermion topological states to a new realm of strongly correlated topological phases whose classifications are much broader Wen2017Zoo . Our prediction broadly expands the basic notion of non-Abelian statistics in combining with symmetry protection and with or without strong correlations. Many future important issues deserve the further in-depth study. For instance, while in the present study we have focused on the PZMs protected by Abelian unitary symmetries, extending the results to the scenarios with non-Abelian symmetries, which can protect coexisting PZMs more than two, may generate fundamentally new type of SPNA statistics due to the nonlinear features in the fractional quasiparticles, possibly achieving the universal quantum computation. On the other hand, extending SPNA statistics to higher-dimensional non-Abelian topological states CChan2017 ; ZCGu2021 may yield the non-Abelian anyons beyond zero dimension.
The roles of symmetries in quantum statistics may revolutionize the basic understanding of quantum many-body physics. The present SPNA statistics for PZMs may be naturally applied to introduce new fractional Abelian statistics by considering symmetries. Generally, when there are multiple locally coexisting fractionalized particles, with or without the symmetry protection, those fractional particles may obey fundamentally different quantum statistics. A novel example may be given for semions. Without symmetry protection, it is well-known that semion pairs obey boson statistics, which was a key motivation to propose the anyon superconductors WilczekBook ; WittenAnyonSC ; FisherAnyonSC ; LaughlinAnyonSC . However, if there is symmetry protection, the semion pairs formed from two symmetry sectors may satisfy the fermion rather than boson statistics. These are highly novel topics for future research.
ACKNOWLEDGMENTS
This work was supported by National Key Research and Development Program of China (2021YFA1400900), the National Natural Science Foundation of China (Grants No. 11825401 and No. 12261160368), the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000), and the Strategic Priority Research Program of the Chinese Academy of Science (Grant No. XDB28000000).
Appendix A Unitary symmetry protection condition
We show that the generic unitary symmetry protection condition for a pair of PZMs can be expressed by the action of the unitary symmetry at the local parity operator. By definition, the most general symmetry protection of these PZMs is defined by $[H_{\text{c}},S]\neq 0$, for $H_{\text{c}}$ in any form, where $S$ is a symmetry of the system. To facilitate the derivaition, we work in the bases of eigenstates of local parity operator $\tilde{Q}_{i}$, where $\tilde{Q}_{i}=U$, $\tilde{\alpha}_{i}=\omega^{(N-1)/2}UV$, $\tilde{\beta}_{i}=V$, and $U|q\rangle=\omega^{q}|q\rangle$, $V|q\rangle=|q-1\rangle$ are the $Z$-gate and $X$-gate (shift matrix) in the $N$-qudit formalism. The local coupling is diagonal in this bases $H_{{\rm c}}={\rm diag}(h_{1},h_{2},\cdots,h_{N})$, where $h_{i}$ is real for hermitianity. We further assume that $S|q\rangle$ is still a local parity eigenstate, $S|q\rangle=e^{i\theta(q^{\prime})}|q^{\prime}\rangle$, which can be regarded as a kind of parity superselection rule. The general symmetry protection condition is equivalent to $[H_{{\rm c}},S]=0\Rightarrow H_{{\rm c}}={\rm const.}$, and we have
$(SH_{{\rm c}})_{ij}=$ $\sum S_{ik}H_{kj}=S_{ij}h_{j},$ (31)
$(H_{{\rm c}}S^{\prime})_{ij}=$ $\sum H_{ik}S_{kj}=h_{i}S_{ij},$ (32)
then an $n$-cycle in $S$ yields $n$ identical $h_{i}$ in $H_{{\rm c}}$ and vice versa. Thus we conclude that the unitary symmetry $S$ must be an $N$-cycle. In this case, there exists a local unitary transformation such that
$GSG^{-1}=e^{i\vartheta/N}V^{p},$ (33)
where $p$ and $N$ are coprime and $G$ commutes with the Hamiltonian of the whole system. The action of $G$ rearranges the eigenstates such that $S|q^{\prime}\rangle=e^{i\theta(q)}|q^{\prime}-p\rangle$ where $\sum\theta(q)=\vartheta$ and gauges the phase $\theta(q)$ by sending $|q^{\prime}\rangle$ to $|q^{\prime\prime}\rangle=e^{-i[\theta(1)+\theta(2)+\cdots+\theta(q^{\prime})-q^{\prime}\vartheta/N]}|q^{\prime}\rangle$, then $S|q^{\prime\prime}\rangle=e^{i\vartheta/N}|q^{\prime\prime}-p\rangle$. Thus we have proved the Eq. (33), which can be rewritten in the operator form
$SQ_{i}S^{-1}=\omega^{p}Q_{i},$ (34)
where $Q_{i}=G^{-1}\tilde{Q}_{i}G$ is a new local parity operator. An example can be given in $\mathbb{Z}_{4}$ case where $S\tilde{Q}_{i}S^{-1}=[-\tilde{Q}_{i}-(1+i)\tilde{Q}_{i}^{2}+i\tilde{Q}_{i}^{3}]/2$. In the eigen-bases of $\tilde{Q}_{i}$, $S|1\rangle=|3\rangle$, $S|3\rangle=|2\rangle$, $S|2\rangle=|4\rangle$, $S|4\rangle=|1\rangle$. The local unitary transformation that swaps $|2\rangle$ and $|3\rangle$ is then $G=1/2+[(1-i)\tilde{Q}_{i}+(1+i)\tilde{Q}_{i}^{3}]/4+S(-1+i\tilde{Q}_{i}-\tilde{Q}_{i}^{2}-i\tilde{Q}_{i}^{3})/4-S^{2}/2+S^{3}(-1-\tilde{Q}_{i}+\tilde{Q}_{i}^{2}-\tilde{Q}_{i}^{3})/4$ and the new local parity operator has the property $SQ_{i}S^{-1}=e^{3\pi i/2}Q_{i}$. The Eq. (34) suggests that the symmetry can advance the local parity by $p$-increments where $p$ and $N$ are coprime. For $N=2$, this means that the symmetry operation switches odd/even fermion parity to even/odd one.
Appendix B Symmetry operation on each of the PZMs
We show in this Appendix how the symmetry acts on the PZMs given that $SQ_{i}S^{-1}=\omega^{p}Q_{i}$. Generically under the unitary symmetry $S$, we have
$S\alpha^{\prime}_{i}S^{-1}=e^{i\theta}\alpha^{\prime}_{i}\sum_{n=1}^{N-1}\lambda_{n}Q_{i}^{n},$ (35)
where $\alpha^{\prime}_{i}=G^{-1}\tilde{\alpha}_{i}G$, and it follows that $S\beta^{\prime}_{i}S^{-1}=e^{i\theta}\omega^{-p}\beta^{\prime}_{i}\sum_{n=1}^{N-1}\lambda_{n}Q_{i}^{n}$ with $\beta^{\prime}_{i}=G^{-1}\tilde{\beta}_{i}G$. The parameters $e^{i\theta}$ and $\lambda_{n}$ are constraint by the symmetry $S$. That is, since $S$ is a unitary operation, the algebraic relations are kept under this operation, namely
11 11 $=\left(S\alpha^{\prime}_{i}S^{-1}\right)^{N}=\left(S\beta^{\prime}_{i}S^{-1}\right)^{N},$ (36)
11 11 $=S{\alpha^{\prime}}_{i}^{\dagger}S^{-1}S\alpha^{\prime}_{i}S^{-1}=S{\beta^{\prime}}_{i}^{\dagger}S^{-1}S\beta^{\prime}_{i}S^{-1}.$ (37)
From Eq. (35) and Eq. (37), we find that $\sum_{n=1}^{N-1}\lambda_{n}Q_{i}^{n}$ is unitary and can be replaced by $e^{i\sum_{n=1}^{N-1}\mu_{n}Q_{i}^{n}}$, where $\mu_{n}^{*}=\mu_{N-n}$. Together with Eq. (36), we have $\theta=2\pi p_{1}/N$, where $p_{1}$ is an integer. Collecting all the facts, we find the action of $S$ on the PZMs can be written as
$S\alpha^{\prime}_{i}S^{-1}=$ $\omega^{p_{1}}\alpha^{\prime}_{i}e^{i\sum_{n=1}^{N-1}\mu_{n}Q_{i}^{n}},$ (38)
$S\beta^{\prime}_{i}S^{-1}=$ $\omega^{p_{2}}\beta^{\prime}_{i}e^{i\sum_{n=1}^{N-1}\mu_{n}Q_{i}^{n}},$ (39)
where $p_{1}-p_{2}=p$, $\mu_{n}^{*}=\mu_{N-n}$. In the case of MZM where $N=2$, these equations are simply $S\alpha^{\prime}_{i}S^{-1}=\pm\alpha^{\prime}_{i}e^{i\mu_{1}Q_{i}}=\pm(\cos\mu_{1}\alpha^{\prime}_{i}+\sin\mu_{1}\beta^{\prime}_{i})$ and $S\beta^{\prime}_{i}S^{-1}=\mp\beta^{\prime}_{i}e^{i\mu_{1}Q_{i}}=\mp(-\sin\mu_{1}\alpha^{\prime}_{i}+\cos\mu_{1}\beta^{\prime}_{i})$, i.e., each MZM is transformed by the symmetry operator into a linear combination of the Majorana doublet.
Appendix C Mixing of two sectors in the new PZMs under perturbation
In this Appendix we show how the specific symmetric perturbation $H_{p}$ in Eq. (29) mix the two sectors. Suppose that $f/J\ll 1$, $V/J\ll 1$ and $f/V\sim\mathcal{O}(1)$, the PZM can be constructed perturbatively as
$\bar{\alpha}_{1}=\alpha_{1}+\sum_{i}m_{i}\hat{O}_{i}^{(1)}+\sum_{j}n_{j}\hat{O}_{j}^{(2)}+\cdots,$ (40)
where $m_{i}\sim\mathcal{O}(f/J)\sim\mathcal{O}(V/J)$ and $n_{j}\sim\mathcal{O}(f^{2}/J^{2})\sim\mathcal{O}(V^{2}/J^{2})\sim\mathcal{O}(fV/J^{2})$. To the zeroth order, we have $\tilde{\alpha}_{1}=\alpha_{1}$ and the following commutator
$[H_{\text{latt}},\alpha_{1}]=(\omega^{*}-1)f\alpha_{2}+(1-\omega^{*})f^{*}\alpha_{2}^{\dagger}\alpha_{1}^{2}.$ (41)
To cancel out the term $(\omega^{*}-1)f\alpha_{2}$, we only have two choices $\alpha_{3}$ and $\alpha_{2}^{2}\alpha_{3}^{\dagger}$ in the first order perturbation, thus at least one of them is involved in $\sum_{i}m_{i}\hat{O}_{i}^{(1)}$. They give the following commutators
$[H_{\text{latt}},\alpha_{3}]$ $=\hat{C}_{1}(\alpha)+\hat{C}_{2}(\alpha,\beta),$ (42)
$[H_{\text{latt}},\alpha_{2}^{2}\alpha_{3}^{\dagger}]$ $=\hat{D}_{1}(\alpha)+\hat{D}_{2}(\alpha)+\hat{D}_{3}(\alpha,\beta),$ (43)
where $\hat{C}_{1}(\alpha)=(1-\omega)J\alpha_{2}^{\dagger}\alpha_{3}^{2}+(\omega-1)J^{*}\alpha_{2}$, $\hat{C}_{2}(\alpha,\beta)=(1-\omega)V\alpha_{2}^{\dagger}\alpha_{3}^{2}\beta_{2}^{\dagger}\beta_{3}+(\omega-1)V^{*}\alpha_{2}\beta_{3}^{\dagger}\beta_{2}$ and $\hat{D}_{1}(\alpha)=(\omega^{-2}-\omega^{*})J\alpha_{2}+(\omega^{-3}-\omega^{-4})J^{*}\alpha_{3}^{-2}\alpha_{2}^{3}$, $\hat{D}_{2}(\alpha)=(1-\omega^{2})f\alpha_{1}^{\dagger}\alpha_{2}^{3}\alpha_{3}^{\dagger}+(1-\omega^{-2})f^{*}\alpha_{3}^{\dagger}\alpha_{2}\alpha_{1}$, $\hat{D}_{3}(\alpha,\beta)=(\omega^{-2}-\omega^{*})V\alpha_{2}\beta_{2}^{\dagger}\beta_{3}+(\omega^{-3}-\omega^{-4})V^{*}\alpha_{3}^{-2}\alpha_{2}^{3}\beta_{3}^{\dagger}\beta_{2}$. For each commuator, there are terms with $\beta_{i}$, then we are left with two possibilities: (i) the terms with $\beta_{i}$ are canceled out in the first order perturbation $[H_{p},\hat{O}_{i}^{(1)}]$, then one can check that certain terms with $\beta_{i}$ like $\alpha_{1}\beta_{3}^{\dagger}\beta_{2}$ or $\alpha_{1}^{\dagger}\alpha_{2}^{2}\beta_{3}^{\dagger}\beta_{2}$ or $\alpha_{2}^{2}\alpha_{3}^{\dagger}\beta_{2}^{2}\beta_{3}^{-2}$ are involved in $\hat{O}_{i}^{(1)}$; (ii) the terms with $\beta_{i}$ are canceled out in the second order perturbation $[H_{\text{latt}}-H_{p},\hat{O}_{j}^{(2)}]$, then certain terms with $\beta_{i}$ like $\alpha_{2}^{2}\alpha_{3}^{\dagger}\beta_{2}\beta_{3}^{\dagger}$ or $\alpha_{3}\beta_{3}^{\dagger}\beta_{2}$ are involved in $\hat{O}_{j}^{(2)}$. In conclusion, $\tilde{\alpha}_{1}$ cannot be constructed solely by terms with $\alpha_{i}$ , and its dynamics is governed by the full Hamiltonian rather than just $H_{\alpha}$. Note that in the above iteration procedure, $\beta_{j}^{\dagger}$ and $\beta_{i}$ always co-occur in $\hat{O}_{i}^{(n)}$, and $S\beta_{j}^{\dagger}\beta_{i}S^{-1}=\omega^{*}\omega\beta_{j}^{\dagger}\beta_{i}=\beta_{j}^{\dagger}\beta_{i}$, we find $\bar{\alpha}_{1}$ still eigenmode of the symmetry $S\bar{\alpha}_{1}S^{-1}=\bar{\alpha}_{1}$.